# Reinforcement Learning for Durable Algorithmic Recourse

### v1

@@ 6 Dicussion and Conlusion
While the RL agent effectively learns environment dynamics, real-world deployment may introduce additional complexities. A key drawback arises during the transient learning phase—especially for the recourse recommender policy—where candidates may receive suboptimal recommendations. This limitation could be mitigated by training this policy on offline data; exploring such an extension constitutes an important step toward practical deployment, and is left for future work.
Moreover, our simulation environment focuses on non-causal recourse, reflecting the work’s emphasis on robustness under competition. The predictor, however, can be paired with any causal recourse method, and the recommender can be trained in environments with underlying causal structures. Additional uncertainties could also be considered, such as candidates unexpectedly not reapplying or shifts in new applicants’ distribution. Exploring these extensions provides a promising path for validating robustness and optimality in more complex, causally grounded settings.
Overall, this work establishes a foundation for durable and adaptive recourse under competition, while opening multiple pathways for further research.
@@ Appendix A Motivating example
To demonstrate the limitations of existing recourse methods, we examine a Ph.D. admission process. Decisions on admissions are determined by an automated system $M(\cdot)$ that evaluates applicants using criteria such as their GPA, educational background, publications, awards, extracurricular activities, English proficiency, and admission test scores. Admission is granted to the top $k$ applicants, where $k$ remains a constant value representing the number of seats available annually.
@@
Finally, recommendations must consider long implementation times, and reapplication delays. Following Venkatasubramanian and Alfano [2020], we argue that recourse should either be permanent or come with an explicit expiration. We adopt the latter and associate each recommendation with a validity horizon $T$, during which the recommendation ensures acceptance. This allows candidates to plan longer-term changes with confidence that their efforts will remain relevant.
@@ Appendix D Experimental Setup
The first step in our experimental setup is to construct the score-based decision model $M(\cdot)$. We generate a synthetic dataset of 10,000 candidates, each described by 10 features and a binary ground-truth label indicating past acceptance or rejection. $M(\cdot)$ is a logistic regression model, trained on this dataset to approximate the ground-truth labels. The model’s probabilistic outputs serve as candidate scores, representing the estimated likelihood of acceptance.
@@ E.1 Recourse Recommender Policy Performance
Method $e_{t}$ $c_{t}$
Ours $1.9\times 10^{-3\phantom{6}}$ $5.9\times 10^{-2}$
Ustun $2.2\times 10^{-16}$ $3.0\times 10^{-1}$
Wachter $2.6\times 10^{-3\phantom{6}}$ $2.7\times 10^{-1}$
DiCE $1.6\times 10^{-2\phantom{6}}$ $3.6\times 10^{-1}$
@@ E.2 Comparison with Wachter and DiCE
Ours $0.707$ $0.365$
Ustun $0.625$ $0.256$
Wachter $0.606$ $0.303$
DiCE $0.550$ $0.247$
Table 3: Recourse feasibility ($\text{RF}_{t}^{T}$), averaged over ten evaluation episodes, for a fixed recourse validity ($\text{RR}_{t}^{T}\approx 0.95$) and $T=1$, across different values of $\beta$.

### v2

@@ 6 Discussion and Conclusions
While the RL agent effectively learns environment dynamics, real-world deployment may introduce additional complexities. A key drawback arises during the transient learning phase—especially for the recourse recommender—where candidates may receive suboptimal recommendations. This limitation could be mitigated by training this policy on offline data, such as data from previous recourse systems used in the same setting, which record how the candidate population evolved in response to recommendations. Exploring such and extension constitutes an important step toward practical deployment and is left for future work.
Moreover, our simulation environment focuses on non-causal recourse, reflecting the work’s emphasis on robustness under competition. The predictor, however, can be paired with any causal recourse method, and the recommender can be trained in environments with underlying causal structures. Additional sources of uncertainty, such as shifts in new applicants’ distribution or in the prediction model, could also be modeled (Appendix B.5). Exploring these extensions provides a promising path for validating robustness in more complex, causally grounded settings.
Lastly, while our framework empirically shows that RL can be leveraged in competitive recourse settings to balance multiple desiderata, no theoretical guarantees of convergence to such a solution can be established, as the chosen RL method (Haarnoja et al., 2018) lacks such guarantees (Appendix H). Overall, this work establishes a foundation for durable and adaptive recourse under competition, while opening multiple pathways for further research.
@@ Appendix A Motivating example
To demonstrate the limitations of existing recourse methods, we examine a Ph.D. admission process. Decisions on admissions are supported by a screening system $M(\cdot)$ that evaluates applicants using criteria such as their GPA, educational background, publications, awards, extracurricular activities, English proficiency, and admission test scores. Admission to the next stage is granted to the top $k$ applicants, where $k$ remains a constant value representing the number of seats available annually.
@@
Finally, recommendations must consider long implementation times and reapplication delays. Following Venkatasubramanian and Alfano [2020], we argue that recourse should either be permanent or come with an explicit expiration date. Offering candidates recommendations that are only valid for a single time step risks creating a false sense of agency, since they may have no realistic way to implement the required changes within that interval. We adopt the latter option and associate each recommendation with a validity horizon $T$, during which the recommendation ensures acceptance. This allows candidates to plan longer-term changes with confidence that their efforts will remain relevant.
In real deployments, practitioners could select $T$ by combining domain knowledge with empirical evidence. Domain expertise can help estimate the typical time required for individuals to implement meaningful changes to key features, allowing stakeholders to choose a horizon that balances recommendation validity with practical feasibility. Additionally, if historical data were available on applicants who received recommendations and later reapplied, one could empirically estimate the distribution of reapplication intervals or the time needed to implement specific changes.
We note that although the responsibility to ensure longer durability may not be as strong as the responsibility to ensure validity (since the latter corresponds more directly to breaking a promise), both are tied to user trust and to the system’s broader accountability. Failures on either front can lead candidates to disregard the recommendations altogether, ultimately rendering the recourse system ineffective.
@@ B.1 Synthetic dataset generation
Our design extends previous simulation environments modeling competitive recourse in limited-resource settings [Fonseca et al., 2023, Bell et al., 2024]. Prior work considered candidates generated in a 2-dimensional feature space, sampled independently at random, where $x=(x_{1},x_{2})$ and $x_{i}\sim\mathcal{N}(\mu=0.5,\sigma=0.3)$ for $i=1,2$. They trained a simple logistic regressor as a classifier, with target variables $y_{i}$ drawn from a binomial distribution.
We improve this synthetic data generation process by considering $10$ continuous features instead of $2$, with each feature sampled from a distinct Gaussian. Moreover, we train the logistic regressor on a dataset constructed in this way, where the ground-truth target is correlated with the features, while still incorporating noise.
@@ B.4 Probability of Reapplying
B.5 Possible Environment Extensions
Our simulation environment is intentionally designed to be extensible, allowing richer behavioral dynamics to be incorporated as needed. In this way, future work could capture more complex candidate behaviors and assess the agent’s ability to learn an effective policy in the presence of additional sources of noise.
Table 1 summarizes several phenomena that could be included, together with indicative implementation strategies.
Table 1: Potential extensions to the simulation environment and corresponding implementation strategies.
Phenomenon Implementation strategy
Heterogeneous urgency and self-confidence Introduce personalized parameters in $p_{\text{reapply}}$, sampled per candidate.
Collective action among candidates Assign a small probability (e.g., $0.03$) to a coordinated non-engagement event. If the event occurs, set $p_{\text{success}}=0$ for every rejected candidate and set $p_{\text{reapply}}=1$ at the next time step.
A priori low trust in the recourse system For each candidate receiving a recommendation, sample a low-probability event (e.g., $0.05$). If the event occurs, set that candidate’s $p_{\text{success}}=0$, regardless of the attainability (for all the time-steps where such candidate reapplies).
Exogenous variation in candidate features Sample a low-probability event (e.g., $0.05$). If the event occurs, select a subset of features and shift the mean or standard deviation of their sampling distributions.
Exogenous shifts in the predictive model With small probability (e.g., $0.05$), retrain the model on a modified dataset where the weights for computing the ground truth are slightly perturbed.
Each phenomenon requires modifications to specific components of the environment. For instance, personalization of candidates’ urgency or self-confidence can be reflected by introducing per-candidate parameters in the reapplication model, scaling the exponent in Equation 9 and $u_{j}$ in Equation 10.
Collective action can be modeled as a rare global event in which candidates strategically choose not to pursue recommendations to avoid intensifying competition. Under such an event, candidates follow their usual dropout dynamics via $p_{\text{dropout}}$, but all remaining candidates reapply at the first available time step with unchanged features (i.e., $p_{\text{success}}=0$ and $p_{\text{reapply}}=1$).
Even when collective action does not occur, candidates may individually decline to engage due to low trust in the recourse system. These candidates similarly keep their features unchanged and reapply at the next opportunity (i.e., $p_{\text{success}}=0$ and $p_{\text{reapply}}=1$), regardless of the attainability of their recommendations.
Exogenous shifts may arise in both the candidate population and the predictive model $M(\cdot)$. Shifts in candidate features may be implemented by perturbing the sampling distributions of selected features, modifying their means or standard deviations. For model shifts, retraining $M(\cdot)$ on a perturbed version of the synthetic training set (whose ground-truth weights are slightly altered) provides a simple mechanism to modify the relationship between features and outcomes, thereby affecting the model’s learned weights.
@@ Appendix D Experimental Setup
The first step in our experimental setup is to construct the score-based decision model $M(\cdot)$. We generate a synthetic dataset of 10 00010\,000 candidates, each described by 10 features and a binary ground-truth label indicating past acceptance or rejection. $M(\cdot)$ is a logistic regression model, trained on this dataset to approximate the ground-truth labels. The model’s probabilistic outputs serve as candidate scores, representing the estimated likelihood of acceptance.
@@ E.1 Recourse Recommender Policy Performance
Method $e_{t}$ $c_{t}$
Ours $1.9\times 10^{-3\phantom{6}}$ $5.9\times 10^{-2}$
Ustun $2.2\times 10^{-16}$ $3.0\times 10^{-1}$
Wachter $2.6\times 10^{-3\phantom{6}}$ $2.7\times 10^{-1}$
DiCE $1.6\times 10^{-2\phantom{6}}$ $3.6\times 10^{-1}$
@@ E.2 Comparison with Wachter and DiCE
Table 4: Recourse feasibility ($\text{RF}_{t}^{T}$), averaged over ten evaluation episodes, for a fixed recourse reliability ($\text{RR}_{t}^{T}\approx 0.95$) and $T=1$, across different values of $\beta$.
@@
Ours $0.71\pm 0.01$ $0.30\pm 0.02$
Ustun $0.63\pm 0.03$ $0.26\pm 0.02$
Wachter $0.61\pm 0.02$ $0.30\pm 0.02$
DiCE $0.55\pm 0.02$ $0.25\pm 0.03$
@@ E.4 Impact of $\beta$ on the Validity-Feasibility Trade-off
Appendix F In-depth comparison of our approach, the hybrid (based on Ustun et al.) and ARR (Dominiguez-Olmedo et al.)
Table 5: Recourse feasibility ($\text{RF}_{t}^{T}$), averaged over ten evaluation episodes, for a fixed recourse validity ($\text{RR}_{t}^{T}\approx 0.95$), varying $T\in\{1,5\}$ and $\beta\in\{0.05,0.01\}$.
$T=1$ $T=5$
Method $\beta=0.05$ $\beta=0.01$ $\beta=0.05$ $\beta=0.01$
Ours $0.71\pm 0.01$ $0.30\pm 0.02$ $0.45\pm 0.01$ $0.25\pm 0.01$
Hybrid (Ustun et al.) $0.63\pm 0.03$ $0.26\pm 0.02$ $0.42\pm 0.03$ $0.22\pm 0.02$
ARR (Dominiguez-Olmedo et al.) $0.57\pm 0.02$ $0.22\pm 0.02$ $0.36\pm 0.02$ $0.16\pm 0.01$
In Table 5, we zoom into the results of our approach, the hybrid variant based on Ustun et al. [2019], and the ARR approach [Dominguez-Olmedo et al., 2022] in high-reliability regimes (i.e., $RR_{t}=0.95$). As shown in the table, our method achieves higher feasibility at the same reliability level than the hybrid approach, by prioritizing changes on features with lower difficulties. Meanwhile, the hybrid approach, which uses our recommender to choose a target score based on the environment’s characteristics and candidates’ behavior, outperforms the ARR method, whose recommendations consist of a target score derived from a robustified threshold (parameterized by $\varepsilon$) at each time step.
The higher feasibility observed in this case is likely due to the fact that, while ARR robustifies the threshold at every time step, our recommender predicts future competition. As a consequence, in some cases it anticipates a decrease in competition in subsequent time steps and therefore lowers the target score.
Specifically, when examining the behavior of our trained recommender $\mu$ (paired with Ustun) in a short evaluation episode (Table 6) with $\beta=0.05$ and $T=1$, we observe that there are time steps in which the agent recommends a target score lower than the most recent threshold ($t=2$ and $t=4$). These recommendations still yield high levels of reliability. As a result, the average recourse feasibility $RF_{t}$ increases, since candidates are only recommended strong changes when necessary.
Table 6: Values of the threshold, the subsequent score recommended by the predictor $\mu$, and the observed Recourse Reliability ($RR_{t}$) and Recourse Feasibility ($RF_{t}$) at each time step in an 8-step evaluation episode. The quantities $RR_{t}$ and $RF_{t}$ are computed relative to the recommendation issued in the previous step. At $t=2$ and $t=4$, the agent recommends a target score lower than the most recently observed threshold. The corresponding values of $RR_{t}$ and $RF_{t}$ at the following steps ($t=3$ and $t=5$) are both desirable.
Time step
$t=1$ $t=2$ $t=3$ $t=4$ $t=5$ $t=6$ $t=7$ $t=8$
Threshold values 0.46 0.63 0.60 0.63 0.56 0.57 0.59 0.54
Recommended scores 0.63 0.60 0.63 0.63 0.58 0.59 0.60
Recourse Reliability $RR_{t}$ 1.0 1.0 1.0 1.0 1.0 1.0 1.0
Recourse Feasibility $RF_{t}$ 0.36 0.67 0.30 0.86 0.57 0.71 0.75
Table 7: Values of the threshold, the subsequent score recommended by ARR, and the observed Recourse Reliability ($RR_{t}$) and Recourse Feasibility ($RF_{t}$) at each time step in an 8-step evaluation episode. The quantities $RR_{t}$ and $RF_{t}$ are computed relative to the recommendation issued in the previous step.
Time step
$t=1$ $t=2$ $t=3$ $t=4$ $t=5$ $t=6$ $t=7$ $t=8$
Threshold values 0.55 0.57 0.59 0.60 0.58 0.62 0.55 0.61
Recommended scores 0.61 0.63 0.65 0.66 0.64 0.68 0.61
Recourse Reliability $RR_{t}$ 1.0 1.0 1.0 1.0 1.0 1.0 1.0
Recourse Feasibility $RF_{t}$ 0.36 0.44 0.33 0.71 0.71 0.43 0.50
Instead, when examining the behavior of ARR in a short evaluation episode (Table 7), we observe that, as expected, the score recommended by ARR is always higher than the most recently observed threshold. As a consequence, the values of recourse feasibility $RF_{t}$ are generally lower.
Appendix G Experiment on German
To assess the suitability of our method in real-world settings, we conducted an additional experiment using the German Credit dataset [^2] https://archive.ics.uci.edu/dataset/522/south+german+credit. Although information on how candidates respond to recommendations is not available in this dataset—a limitation relative to literature datasets noted by prior work on recourse in competitive settings [Fonseca et al., 2023]—it can still be used to sample candidates’ initial features, grounding them in a realistic setting. Below, we describe how our experimental setup is adapted for this dataset and present preliminary results.
G.1 Experimental set-up
We use the German Credit dataset to train the predictive model $M(\cdot)$ and to sample candidates’ initial features. Since the dataset only contains 1 0001\,000 samples, which is insufficient to both train $M(\cdot)$ and generate a diverse set of candidate profiles for all time steps across episodes, we augment it with synthetic data, generated to be as similar as possible to the original samples.
To preserve the statistical properties of the original dataset, we employ a Gaussian copula approach [Nelsen, 2006], that captures both the marginal distributions and the correlation structure among features, thus ensuring that synthetic samples closely resemble the original data in terms of both individual feature distributions and inter-feature dependencies.
A key difference between the German dataset and the synthetic dataset that we use in our main experiments is the presence of categorical features. This changes the meaning of feature difficulties: while for continuous features they represent a relative difficulty to implement a feature change, in the case of categorical features they represent the difficulty of switching from a category to another. Thus, for every categorical feature we define a matrix of difficulty parameters, one for every ordered pair of categories. These parameters vary both within the same category (changing the savings from <100<100 to $\geq 1000$ is more difficult than changing it to $100\leq x<1000$), and among different categories (changing the purpose of the loan is easier than changing the savings or the employment). These values represent the difficulty of implementing the corresponding category switch.
The probability of successfully implementing a categorical feature switch $p_{\text{success}}$ is also different from Equation 8, due to these changes. Specifically, we design it as:
$p_{\text{success}}^{\text{cat}}=(1-d_{i,j})^{\frac{1}{\beta_{\text{cat}}}}$ (19)
where $\beta_{\text{cat}}$ is a setting difficulty parameter, analogous to the parameter $\beta$ previously defined. Lower values of $\beta_{\text{cat}}$ indicate settings where implementing changes to categorical features is generally more difficult. Instead, $d_{i,j}$ represents the difficulty of switching from category $i$ to category $j$.
G.2 Implementation details
We train our recommender agent in an environment similar to the one described in Section 3.1 and Appendix B, with two key modifications: (i) recommendations on categorical features are constrained to category switches, and (ii) the probability of successfully implementing a category switch follows the formulation defined in Equation 19.
The recommender requires a longer training period compared to the synthetic dataset experiments. This increased complexity likely stems from the challenges inherent in categorical feature recommendations. Specifically, with 17 categorical features spanning 54 distinct categories, the agent must navigate a substantially larger action space of possible feature switches. Moreover, category switches induce non-smooth changes in the credit score, making it difficult to identify counterfactuals that achieve a precise target score. To account for this complexity, we train the recommender for 20 00020\,000 time-steps in the first phase and 30 00030\,000 time-steps in the second phase.
The predictor, as in the experiments on synthetic samples, is trained for 7 0007\,000 time-steps. We use the following hyperparameters: $T=1$, $\beta=0.05$ and $\beta_{\text{cat}}=1.0$.
G.3 Results
In Table 8, we report the results of the preliminary experiment on the German dataset. As can be seen, both our method and the hybrid variant achieve desirable levels of reliability in this more complex setting. On the other hand, as in the scenario with synthetic data, the baseline’s threshold-based policy yields recommendations that too many candidates can satisfy, thereby reducing reliability.
These results showcase the utility of our approach in more complex scenarios, where feature values are drawn from a real dataset and both continuous and categorical features are taken into account.
Table 8: Results for the baseline, the hybrid method (based on Ustun), and our approach in the simulation environment based on German, with $\beta=0.05$, $\beta_{\text{cat}}=1.0$, and $T=1$.
Method
Baseline Hybrid Ours
Recourse Reliability $RR_{t}$ $0.43\pm 0.05$ $0.90\pm 0.04$ $0.89\pm 0.03$
Recourse Feasibility $RF_{t}$ $0.78\pm 0.06$ $0.36\pm 0.05$ $0.40\pm 0.03$
Appendix H Discussion on convergence guarantees
While our method and the hybrid variants yield desirable solutions in all four settings under consideration ($T\in\{1,5\}$, $\beta\in\{0.05,0.01\}$), we cannot provide theoretical guarantees that the agent converges to an optimal solution. This limitation arises because the reinforcement learning algorithm we employ, Soft Actor Critic (SAC) [Haarnoja et al., 2018], does not include such guarantees.
In their paper, the authors provide lemmas and theorems establishing convergence of Soft Policy Iteration, the tabular algorithm SAC is based on. However, these results rely on assumptions that SAC does not satisfy. SAC uses neural networks to approximate policies and value functions, which breaks the assumption of exact or tabular representations, and it operates in continuous action spaces. Moreover, SAC incorporates entropy maximization and stochastic actor updates, which further depart from the theoretical setting in which convergence can be proven.
At the same time, deriving strong theoretical bounds in our environment would be extremely difficult, due to the complexity of the environment. Candidates’ behavior is noisy, the system is only partially observable, and the dynamics involve significant stochasticity. Given these complexities, this work is focused on establishing a rigorous empirical framework that demonstrates the practical effectiveness of our approach across diverse experimental conditions.
Appendix I Use of Large Language Models
Large language models were used solely to improve the clarity and grammar of the text and to generate the icons of candidates in Figure 1. All substantive content was written by the authors; LLMs were applied only for minor phrasing refinements.